OpenStudy (anonymous):

for the given polynomial P(x) and the given c, use the remainder theorem to fid P(c). P(x)=x^3+3x^2-7x=4 ; c=1

OpenStudy (anonymous):

P(x)=4? you sure thats supposed to = 4?

OpenStudy (anonymous):

P(x)=x^3+3x^2-7x+4 ; c=1

OpenStudy (campbell_st):

find P(1) = (1)^3 + 3(1)^2 - 7(1) + 4 P(1) = -1 then when x^3 + 3x^2 -7x +4 is divided by (x -1) the remainder is -1

OpenStudy (anonymous):

1 3 -7 4 0 1 4 -3 1 4 -3 1 therefore P(1) is 1 for P(1)